A stock just paid a dividend of $0.90. Its dividends are expected to grow at a rate of 8.50% for the next two years, at 7.50% for the next year after that, and then adjust down to a long-term growth rate of 2.00%. If your required return is 11.50%, then what is your estimate of this stock’s current value? Give your response to two decimal places.
Answer #1

D1=(0.9*1.085)=0.9765

D2=(0.9765*1.085)=1.0595025

D3=(1.0595025*1.075)=1.138965188

Value after year 3=(D3*Growth Rate)/(Required rate-Growth rate)

=(1.138965188*1.02)/(0.115-0.02)

=$12.22888939

Hence current price=Future dividends and value*Present value of discounting factor(rate%,time period)

=0.9765/1.115+1.0595025/1.115^2+1.138965188/1.115^3+$12.22888939/1.115^3

=$11.37(Approx).